OpenStudy (anonymous):

Edison has a dog walking business. He earns $6.00 for each dog that he walks. Edison walked 4 dogs on Friday, 6 dogs on Saturday, and some additional dogs on Sunday. He earned a total of $108.00 for those three days. How many dogs did he walk on Sunday? A. 8 B. 10 C. 9 D. 6

OpenStudy (anonymous):

4(6) +6(6) + x(6)=108 24+36+x(6)=108 60+x(6)+108 -60 -60 x(6)=48 (Divide by 6 both sides) x=8

OpenStudy (anonymous):

4(6) +6(6) + 6x=108 24+36+6x=108 60+6x+108 -60 -60 6x=48 x=8 Less confusing